Mags Ltd. has purchased a new machine for a cost of $500,000. The machine has just been

installed and the cost of installation is $30,000. The internal auditors have advised that the

cost of installation cannot be depreciated. The machine’s suppliers have requested a 20%

deposit on installation with the remainder to be paid within six months. The current

estimated before-tax net operating cash revenue for the coming four years are $300,000 in

the first year, $320,000 in the second year, $340,000 in the third year and $360,000 in the

fourth year. The new machine is expected to increase the expected before-tax net operating

cash revenue for the next four years by 70% of the current estimated value. Mags will take

out a business loan to fund the purchase of the machine and the interest payments on the

loan will be $100,000 per annum. The machine will be sold at the end of the fourth year and

its estimated market value at that time is $65,000. The company tax rate is 30% and

reducing balance depreciation is permitted. The required rate of return is 13% per year and

the prevailing market interest rate is 6% per year.

Prepare a cash flow analysis for the useful economic life of the machine, and use the cash

flow analysis to estimate the machine’s net present value. Show all workings.
Answer #1
CALCULATION OF DEPRECIATION TAX SHIELD
Depreciation Rate=r=1-((S/C)^(1/n))
S=Residual value=$65000
C=Original Cost=$500000
n= Life =4 years
r=1-((65000/500000)^(1/4))=1-0.600462 1.0000000
r=0.3995376
A B=A*0.3995376 C=A-B D=B*30%
Year Beginning book Value Depreiation Amount Ending Book Value Depreciation Tax Shield
1 $500,000 $199,769 $300,231 $59,931
2 $300,231 $119,954 $180,278 $35,986
3 $180,278 $72,028 $108,250 $21,608
4 $108,250 $43,250 $65,000 $12,975
Annual Interest Tax Shield=100000*30% $30,000
Payment for installation in year 0=30000*20% $6,000
Balance Payment for installation at end of year=0.5 $24,000 (30000-6000)
Present Value of balance Payment =24000/(1.13^0.5) $22,577
Present Value of initial Cash Flow =6000+22577= $28,577
N Year 0 1 2 3 4
A Present Value of initial Cash out Flow ($28,577)
B Interest Payment ($100,000) ($100,000) ($100,000) ($100,000)
C Principal Payment ($500,000)
D Salvage Value cash flow $65,000
E Before tax net operating cash revenue $300,000 $320,000 $340,000 $360,000
F=E*70% Increase in before tax revenue $210,000 $224,000 $238,000 $252,000
G=F*(1-0.3) Increase in After tax revenue $147,000 $156,800 $166,600 $176,400
H Depreciation tax shield $59,931 $35,986 $21,608 $12,975
I Interest Tax Shield $30,000 $30,000 $30,000 $30,000
J=G+H+I Net Increase in annual operating cash flows $236,931 $222,786 $218,208 $219,375
K=A+B+C+D+J NET CASH FLOW ($28,577) $136,931 $122,786 $118,208 ($315,625) SUM
PV=K/(1.13^N) Present Value of Net Cash Flow ($28,577) $121,178 $96,160 $81,924 -$193,579 $77,105
NPV=sum of PV NET PRESENT VALUE (NPV) $77,105